OpenStudy (some.random.cool.kid):

Solve the system y equals negative x squared, y equals x minus two.

OpenStudy (some.random.cool.kid):

@ganeshie8 @Zarkon @zzr0ck3r @Nnesha

OpenStudy (some.random.cool.kid):

@undeadknight26 @peachpi

OpenStudy (some.random.cool.kid):

@ParthKohli

Nnesha (nnesha):

\[\huge\rm y=-x^2\]\[\huge\rm \color{ReD}{y}=x-2\] substitute red y for first equation

Nnesha (nnesha):

 and then solve the quadratic equation